Joe is 1.6 m tall. His shadow is 2m long when he stands 3m from the base of a floodlight. What is the height of the floodlight?
Answer by checkley77(12844)

I don't understand your question regarding where I got the 6.
There is no 6 in this problem other than the 1.6 meters for Joe's height.
-------------------------------------------------------------------------
1.6/2=x/(2+3)
1.6/2=x/5
2x=5*1.6
2x=8
x=8/2
x=4 feet is the height of the floodlight.